# X-ray panorama of the SS 433/W50 complex by SRG/eROSITA

### v1

@@
Galactic microquasar SS 433 and the radio nebula W50 surrounding it present a prototypical example of a hyper-Eddington binary system shaping its ambient interstellar medium via energetic outflows. In this paper, we present X-ray observations of the SS 433/W50 complex by the eROSITA telescope onboard the SRG space observatory. These data provide images of the entire nebula characterized by a very large dynamic range and allow spectral analysis of the diffuse X-ray emission. In particular, these data illustrate a close connection between the thermal and non-thermal components of W50 on scales ranging from sub-parsecs, represented by narrow X-ray bright filaments, to the entire extent of the nebula. These data also allow us to fully characterize a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts, but are prominent in very high energy gamma-ray emission. The resulting multifaceted picture of the interaction between energetic outflows and the surrounding medium paves the way for future focused multiwavelength observations and dedicated numerical simulations.
Key Words.:
ISM: jets and outflows – X-rays: binaries – plasmas – acceleration of particles1 Introduction
@@
SS 433 is a unique close stellar binary system in our Galaxy (located at and kpc) composed of a relativistic compact object (a black hole or a neutron star) which accretes matter from its companion at sustainably hyper-Eddington rate (see Fabrika, 2004; Cherepashchuk et al., 2025, for reviews). Having been first listed in the catalog of point-like H emitters (Stephenson & Sanduleak, 1977), it quickly became one of the most studied astrophysical objects thanks to its very peculiar optical spectrum (Margon et al., 1979) featuring pairs of narrow (with half-opening angle ) blue- and red-shifted lines (with respect to their rest frame positions) with the shifts varying periodically in time (see Margon, 1984, for a review of early observational results). A kinematic model associates this emission with a pair of narrow relativistic jets having constant bulk velocity of a quarter of the speed of light and changing direction in a regular precession manner with a period days (Abell & Margon, 1979; Fabian & Rees, 1979). This period is much longer than the period of orbital modulation ( days) and is interpreted as due to precession of a thick accretion disk with the amplitude of 21∘ and the mean axis inclined by the angle to the line of sight (e.g., Katz, 1980; Sarazin et al., 1980; Whitmire & Matese, 1980). Subsequent detection of extended and variable radio emission (Spencer, 1979) perfectly described by the precessing jets pattern fully confirmed the picture of twin relativistic outflows (Hjellming & Johnston, 1981a, b; Blundell & Bowler, 2004). Moreover, the proper motion of the individual radio-emitting blobs allowed one to constrain the distance to the source with a remarkable 10% accuracy, being kpc (Hjellming & Johnston, 1981a; Blundell & Bowler, 2004; Marshall et al., 2013).
@@
Despite indications of both quasi-regular (nutation, Katz et al., 1982) and sporadic (jittering, Kubota et al., 2010) deviations from the predictions of this simple model as well as presence of (relatively short) periods of time when the Doppler shifts of optical emission lines change significantly (e.g., Blundell et al., 2011; Medvedev et al., 2022), all the derived parameters of the model stay remarkably stable over more than 40 years of continuous monitoring observations (Cherepashchuk et al., 2018). This stability likely reflects extreme robustness of the mass transfer occurring in the hyper-Eddington regime (van den Heuvel, 1981; van den Heuvel et al., 2017) with the estimated accretion rate (e.g., Shklovskii, 1981; Cherepashchuk, 1981; Leibowitz, 1984; Fuchs et al., 2006), where is the mass of the compact object and is the Solar mass and is the critical Eddington mass accretion rate. The compact object is most likely a black hole with (Cherepashchuk et al., 2023), although the possibility of it being a neutron star () has also been discussed (e.g. Goranskij, 2011; Medvedev et al., 2018).
Since X-ray emission coming from the innermost parts of the hyper-Eddington accretion disk is expected to be strongly beamed in the axial direction (Shakura & Sunyaev, 1973; Begelman et al., 2006; Poutanen et al., 2007), we are not able to observe it directly, given the high inclination angle of the system. Indeed, although X-ray emission from this source has also been detected early on (Seward et al., 1976; Forman et al., 1978), its luminosity is relatively low compared to the Eddington luminosity (Grindlay et al., 1984). As a result, SS 433 has long been considered as a beamed away prototype of Ultra Luminous X-ray sources (ULXs, e.g., Kaaret et al., 2017, for a review) observed in normal star-forming galaxies (Fabrika & Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007), although possible indications of the ‘hidden’ central emission are yet to be confirmed (Medvedev & Fabrika, 2010; Khabibullin & Sazonov, 2016, 2019; Middleton et al., 2021; Fogantini et al., 2023).
@@
Instead, what dominates X-ray emission of the system is again thermal emission from a twin pair of the baryonic jets having almost identical orientation and bulk velocity to the optically emitting jets (Watson et al., 1986). Numerous pairs of blue- and red-shifted narrow emission lines of highly ionized heavy elements (primarily, silicon, sulfur, iron and nickel) allow one to build a multi-temperature model of the ballistically expanding and cooling flow (Watson et al., 1986; Kotani et al., 1996; Marshall et al., 2002; Brinkmann et al., 2005; Medvedev & Fabrika, 2010; Marshall et al., 2013; Khabibullin et al., 2016) which first becomes visible when its temperature is above 20 keV and then cools down below 1 keV and likely fragments later on due to thermal instability (Brinkmann et al., 1988). Detailed spectroscopic modeling of this emission (e.g. Khabibullin et al., 2016; Medvedev et al., 2019) in combination with the eclipsing observations of the jets by the companion star (Stewart et al., 1987; Kawai et al., 1989; Filippova et al., 2006; Lopez et al., 2006; Atapin & Fabrika, 2016) enable very robust determination of the jets’ physical properties (so-called jets ‘biometry’, consisting of their size, opening angle, velocity, temperature, and the mass flux). Thanks to this we know very robustly the kinetic luminosity of these outflows, which exceeds (e.g., Medvedev et al., 2019), confirming the necessity of a hyper-Eddington ‘engine’ releasing large amount of energy in the form of relativistic outflows, consistently with the early theoretical (Shakura & Sunyaev, 1973) and modern numerical (e.g., Ohsuga & Mineshige, 2011; Jiang et al., 2019; Yoshioka et al., 2022; Fragile et al., 2025) predictions.
@@
In contrast to the radiation freely escaping the system, the kinetic energy of the launched outflows cannot be easily transported away from the source and has to be deposited into the surrounding interstellar medium. The estimated total amount of this energy exceeds erg for a life-time of the system at the level of 30,000 yr (e.g., van den Heuvel et al., 1980). This energy is at least on par with the energy of the initial supernova explosion that should have accompanied the birth of the compact object, meaning that a quite large region pc in its vicinity can be shaped by this later activity (e.g., Begelman et al., 1980; Zealey et al., 1980; van den Heuvel, 1981; Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Churazov et al., 2020, 2024).
@@
Indeed, SS 433 is known to be surrounded by the famous giant ( pc) radio nebula W50 (Westerhout, 1958; Holden & Caswell, 1969; Geldzahler et al., 1980; Elston & Baum, 1987; Dubner et al., 1998; Goodall et al., 2011b; Broderick et al., 2018; Sakemi et al., 2021), which has a suggestive elongated morphology, with the elongation axis being co-aligned with the symmetry axis of the jets precession cone. The asymmetry of the lobes, or “ears”, of the nebula can be readily explained by the noticeable stratification of the surrounding medium, given that the system is located at the height pc below the Galactic plane and has the size comparable to the scale height of the exponential cold gas disk (e.g., Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a). The interior of the nebula is also known to shine in soft thermal X-rays (Brinkmann et al., 1996, 2007; Safi-Harb & Ögelman, 1997; Chi et al., 2024), hard non-thermal (Seward et al., 1980; Watson et al., 1983; Yamauchi et al., 1994; Safi-Harb & Petre, 1999; Namiki et al., 2000; Safi-Harb et al., 2022; Kaaret et al., 2024) and filamentary H emission (Kirshner & Chevalier, 1980; Shuder et al., 1980; Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017; Rosado et al., 2021), indicative of the high velocity shocks propagating through a relatively dense interstellar medium.
@@
The nature of the extended X-ray jets (EXJ) remains more puzzling, although their morphology, spectra, and polarization all point to the synchrotron origin of the emission (e.g., Safi-Harb et al., 2022; Kaaret et al., 2024). Most recently, Very High Energy (VHE) emission from these structures has been detected and mapped (e.g., Abeysekara et al., 2018; Alfaro et al., 2024; H. E. S. S. Collaboration et al., 2024; LHAASO Collaboration, 2024), demonstrating that indeed ultra-relativistic electrons with energy above 100 TeV are accelerated in the system (e.g, Kimura et al., 2020), probably as a consequence of internal shocks and recollimation in the energetic trans-relativistic outflows (Bykov et al., 2025). The overall production of cosmic rays by such systems (super-Eddington accretors with powerful outflows) might be of importance for the total CR budget and local variations in its spectrum from place to place (e.g., Heinz & Sunyaev, 2002), in particular in the PeV regime (e.g., Peretti et al., 2025; Wang et al., 2025; Bykov et al., 2025).
@@
Although the paradigm of “a super-Eddington X-ray binary powering a nebula with energetic outflows” appears to be well established and consistent with the most salient observational properties of the SS 433/W50 complex, the details of almost all aspects of this interaction remain unclear. Indeed, all the phenomena directly associated with the central source can be trace only up to a distance of mere 0.1 pc from it, where the last signatures of the corkscrew pattern of the precessing jets are observed in radio and X-rays (Migliari et al., 2005; Miller-Jones et al., 2008; Khabibullin & Sazonov, 2017; Sakai et al., 2025). A model invoking steady deceleration and disruption of the jets can, in principle, account for such behavior (Panferov, 2014, 2017), but then the re-appearance of the “jets” with similar to the injection velocities tens of pc away from the central source might be problematic. Morphology of the soft X-ray emission, best mapped by the ROSAT mosaic observations (Brinkmann et al., 1996, 2005), is also quite puzzling, with large global brightness variations and thin filamentary structures, not very commonly observed in similar situations of supernova remnants (which typically feature shell-, plerion- or mixed types of morphology). Deep pointed observations with XMM-Newton, Chandra, Suzaku, and NuSTAR were capable of covering certain parts of the nebula (e.g., Safi-Harb et al., 2022; Chi et al., 2024; Tsuji et al., 2025), but a uniform and sensitive coverage of the entire nebula with adequate spectroscopic capabilities is missing.
@@
Here we present results of X-ray observations of the SS 433/W50 complex by the eROSITA telescope (Predehl et al., 2021) onboard the SRG space observatory (Sunyaev et al., 2021), which were performed during its Performance Verification (PV) phase. The obtained mosaic provides us with X-ray images of the entire extent of W50 with unprecedented dynamic range, bringing X-ray data on par with the radio and other multi-wavelength data, and, in addition, delivers exquisite spectral information for the rich variety of diffuse structures. These data, in particular, demonstrate a close connection between the thermal and non-thermal components of W50 on scales ranging from subparsec, represented by narrow X-ray and radio bright filaments, to the entire extent of the nebula. These data also enable full characterization of a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts.
@@
The structure of the paper is following: we describe the observations and the obtained data in Sect. 2; results of the broad-band and spectrally-resolved imaging are presented in Sect. 3 and Sect. 4, respectively; spectroscopic analysis of the emission from several selected regions follows in Sect. 5; the obtained global picture is discussed in Sect. 6 and summarized in Sect. 7.
@@ 2 Data
Since W50 spans approximately degrees on the sky, corresponding to the physical size of at the distance of 5 kpc111The most recent parallax measurement for SS 433 in Gaia DR3 corresponds to the significantly larger distance of kpc (Gaia DR3 4293406612283985024, Gaia Collaboration et al., 2023), we caution of using this value given large residuals of the astrometric solution. The previous parallax measurement published in Gaia DR2 corresponded to a much smaller distance of kpc (Gaia DR2 4293406612283985024, Arnason et al., 2021)., the observing campaign consisted of 9 individual telescope pointings (each covering a circular area of in diameter). These pointings were arranged in a mosaic to result in a nearly uniform coverage of the entire nebula and, in addition, the highest sensitivity in the most interesting regions. The observing dates, from October 5th to October 8th and then from October 19th to October 20th, 2019, were chosen to coincide with the two consecutive orbital eclipses of the central source SS 433 by its normal companion star (according to orbital ephemerids by Eikenberry et al., 2001). During eclipses, the apparent X-ray flux of SS 433 goes down by a factor of several (e.g., Kawai et al., 1989; Lopez et al., 2006; Medvedev et al., 2019), minimizing possible adverse effects of contamination and pile-up due to the presence of a very bright object in the field-of-view, even though relative suppression of the soft X-ray lines being rather minor (Marshall et al., 2013). The typical effective (i.e., equivalent to the nominal eROSITA on-axis performance) exposure time across the map amounts to 30-50 ks, leading to a factor of deeper images compared to the previously available full-extent maps of W50 by observatory in the 0.5-2 keV band (Brinkmann et al., 1996) and Einstein observatory in the 2-4 keV band (Watson et al., 1983). In addition to that, the eROSITA data have much better spatial ( arcsec on the telescope axis and averaged across the field-of-view) and spectral ( eV at 1 keV) resolution (Predehl et al., 2021). A comparable quality of the data is provided by multiple Chandra and XMM-Newton observations, but their coverage of the whole W50 nebula is non-uniform and still incomplete after more than 20 years of observations (Brinkmann et al., 2005; Safi-Harb et al., 2022; Chi et al., 2024). On the other hand, the SRG/eROSITA data are complementary to these observations, as well as to the imaging observations in harder X-rays with NuSTAR, which offer a unique view of non-thermally emitting populations. In this paper, we focus on the SRG/eROSITA X-ray data alone and leave combined analyses for future studies.
@@
We also supplement the PV data with the data obtained in the course of the SRG/eROSITA All-Sky Survey (Predehl et al., 2021; Sunyaev et al., 2021). Although the survey data are much shallower in sensitivity (the typical exposure time per point is s after combining the data of four scans), they provide uniform coverage of the entire area surrounding the nebula. For them, the standard data reduction, background estimation and exposure- and vignetting corrections were applied in the same way as was done for mapping diffuse emission from nearby galaxy clusters (Churazov et al., 2021a, 2023), supernova remnants (Churazov et al., 2021b; Khabibullin et al., 2023, 2024), and the Galactic diffuse emission (Predehl et al., 2020; Khabibullin et al., 2022; Churazov et al., 2024) on even larger scales. Here we also go one step further and combine the PV and all-sky data in a seamless fashion, similarly as it has been done for Chandra and SRG/eROSITA observations of the Perseus cluster (Churazov et al., 2025).
@@
A comparison of the degrees field in radio and X-ray bands is shown in Fig. 1. Although in the radio image, the outer boundary of the W50 nebula appears to be very prominent, the X-ray image is instead dominated by the very bright central object (SS 433) and elongated structures known as “extended X-ray jets” (EXJ) located well inside the radio boundary of W50 (Seward et al., 1980; Watson et al., 1983; Brinkmann et al., 1996). At a distance of 5 kpc, these images span 350 pc and range from 17 pc to almost 370 pc in distance from the Galactic plane. Given that the scale heights for the cold and warm components of the interstellar medium are tens and hundreds of pc, respectively, one could naturally expect signatures of the environment stratification (e.g., Goodall et al., 2011a). This is indeed revealed by the asymmetry of the W50 lobes, with the one directed towards the Galactic plane (the Western Lobe in equatorial coordinates) being more compact than the opposite one (the Eastern lobe).
@@
On larger scales, clear variations of the background and foreground diffuse X-ray emission are seen. Inspection of 3D reddening maps (e.g., Green et al., 2019) shows a clear anti-correlation of the reddening and soft X-ray surface brightness (see Appendix C), suggesting that absorption of X-rays is the main reason for these variations. Morphologically similar structures appear in the 3D reddening maps at a distance of 2-3 kpc, i.e., in the foreground to W50/SS 433. Therefore, it is plausible that the majority of them are physically unrelated to W50.

### v2

@@
The Galactic microquasar SS 433 and the radio nebula W50 surrounding it present a prototypical example of a hyper-Eddington binary system shaping its ambient interstellar medium via energetic outflows. In this paper, we present X-ray observations of the SS 433/W50 complex by the eROSITA telescope onboard the SRG space observatory. These data provide images of the entire nebula characterized by a very large dynamic range and allow spectral analysis of the diffuse X-ray emission. In particular, these data illustrate a close connection between the thermal and non-thermal components of W50 on scales ranging from sub-parsecs, represented by narrow X-ray bright filaments, to the entire extent of of the nebula. These data also allowed us to fully characterize a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts but are prominent in very-high-energy gamma-ray emission. The resulting multifaceted picture of the interaction between energetic outflows and the surrounding medium paves the way for future focused multiwavelength observations and dedicated numerical simulations.
1 Introduction
@@
SS 433 is a unique close stellar binary system in our Galaxy (located at and kpc) composed of a relativistic compact object (a black hole or a neutron star), which sustainably accretes matter from its companion at a hyper-Eddington rate (see Fabrika, 2004; Cherepashchuk et al., 2025, for reviews). Having first been listed in the catalog of point-like H emitters (Stephenson and Sanduleak, 1977), it quickly became one of the most studied astrophysical objects thanks to its very peculiar optical spectrum (Margon et al., 1979) featuring pairs of narrow (with a half-opening angle of ) blue- and redshifted lines (with respect to their rest-frame positions), with the shifts varying periodically in time (for a review of early observational results see Margon, 1984). A kinematic model associates this emission with a pair of narrow relativistic jets with a constant bulk velocity of a quarter of the speed of light (c) that change direction in a regular precession manner with a period of days (Abell and Margon, 1979; Fabian and Rees, 1979). This period is much longer than the period of orbital modulation ( days) and is interpreted as being due to the precession of a thick accretion disk with an amplitude of 21∘ and mean axis inclined by the angle to the line of sight (e.g., Katz, 1980; Sarazin et al., 1980; Whitmire and Matese, 1980). Subsequent detection of extended and variable radio emission (Spencer, 1979) perfectly described by the precessing jets pattern fully confirmed the picture of twin relativistic outflows (Hjellming and Johnston, 1981a, b; Blundell and Bowler, 2004). Moreover, the proper motion of the individual radio-emitting blobs allowed one to constrain the distance to the source with a remarkable 10% accuracy; i.e., kpc (Hjellming and Johnston, 1981a; Blundell and Bowler, 2004; Marshall et al., 2013).
@@
Despite indications of both quasi-regular (nutation, Katz et al., 1982) and sporadic (jittering, Kubota et al., 2010) deviations from the predictions of this simple model as well as the presence of (relatively short) periods of time when the Doppler shifts of optical emission lines change significantly (e.g., Blundell et al., 2011; Medvedev et al., 2022), all the derived parameters of the model stay remarkably stable over more than 40 years of continuous monitoring observations (Cherepashchuk et al., 2018). This stability likely reflects extreme robustness of the mass transfer occurring in the hyper-Eddington regime (van den Heuvel, 1981; van den Heuvel et al., 2017) with the estimated accretion rate, (e.g., Shklovskii, 1981; Cherepashchuk, 1981; Leibowitz, 1984; Fuchs et al., 2006), where is the mass of the compact object, is the solar mass, and is the critical Eddington mass-accretion rate. The compact object is most likely a black hole with (Cherepashchuk et al., 2023), although the possibility of it being a neutron star () has also been discussed (e.g., Goranskij, 2011; Medvedev et al., 2018).
@@
Since X-ray emission coming from the innermost parts of the hyper-Eddington accretion disk is expected to be strongly beamed in the axial direction (Shakura and Sunyaev, 1973; Begelman et al., 2006; Poutanen et al., 2007), we are not able to observe it directly, given the high inclination angle of the system. Although X-ray emission from this source was detected early on (Seward et al., 1976; Forman et al., 1978), it was also found (e.g., Grindlay et al., 1984) that its apparent luminosity, , is very low compared to the Eddington luminosity, erg s-1, of an compact object. As a result, SS 433 has long been considered as a beamed away prototype of ultra-luminous X-ray sources (ULXs; for a review see Kaaret et al., 2017) observed in normal star-forming galaxies (Fabrika and Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007), although possible indications of the ”hidden” central emission are yet to be confirmed (Medvedev and Fabrika, 2010; Khabibullin and Sazonov, 2016, 2019; Middleton et al., 2021; Fogantini et al., 2023). Instead, what dominates the X-ray emission of the system is again thermal emission from a twin pair of the baryonic jets having almost identical orientation and bulk velocity to the optically emitting jets (Watson et al., 1986). Numerous pairs of blue- and redshifted narrow emission lines of highly ionized heavy elements (primarily silicon, sulfur, iron, and nickel) allow one to build a multi-temperature model of the ballistically expanding and cooling flow (Watson et al., 1986; Kotani et al., 1996; Marshall et al., 2002; Brinkmann et al., 2005; Medvedev and Fabrika, 2010; Marshall et al., 2013; Khabibullin et al., 2016), which first becomes visible when its temperature is above 20 kiloelectronvolt (keV), then cools down below 1 keV, and likely fragments later on due to thermal instability (Brinkmann et al., 1988). Detailed spectroscopic modeling of this emission (e.g., Khabibullin et al., 2016; Medvedev et al., 2019) in combination with the eclipsing observations of the jets by the companion star (Stewart et al., 1987; Kawai et al., 1989; Filippova et al., 2006; Lopez et al., 2006; Atapin and Fabrika, 2016) enable very robust determination of the jets’ physical properties (their so-called ”biometry”, consisting of their size, opening angle, velocity, temperature, and the mass flux). Thanks to this we know the kinetic luminosity of these outflows with certainty; it exceeds (e.g., Medvedev et al., 2019), confirming the necessity of a hyper-Eddington ”engine” releasing large amount of energy in the form of relativistic outflows, consistently with the early theoretical (Shakura and Sunyaev, 1973) and modern numerical (e.g., Ohsuga and Mineshige, 2011; Jiang et al., 2019; Yoshioka et al., 2022; Fragile et al., 2025) predictions.
@@
In contrast to the radiation freely escaping the system, the kinetic energy of the launched outflows cannot be easily transported away from the source and has to be deposited into the surrounding interstellar medium. The estimated total amount of this energy exceeds erg for a lifetime of the system at the level of 30,000 yr (e.g., van den Heuvel et al., 1980). This energy is at least on par with the energy of the initial supernova explosion that accompanied the birth of the compact object, meaning that quite a large region of pc in its vicinity can be shaped by this later activity (e.g., Begelman et al., 1980; Zealey et al., 1980; van den Heuvel, 1981; Velázquez and Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Churazov et al., 2020, 2024).
@@
Indeed, SS 433 is known to be surrounded by the famous giant ( pc) radio nebula W50 (Westerhout, 1958; Holden and Caswell, 1969; Geldzahler et al., 1980; Elston and Baum, 1987; Dubner et al., 1998; Goodall et al., 2011b; Broderick et al., 2018; Sakemi et al., 2021); the latter has a suggestive, elongated morphology, with the elongation axis being co-aligned with the symmetry axis of the jets’ precession cone. The asymmetry of the lobes, or “ears”, of the nebula can be readily explained by the noticeable stratification of the surrounding medium, given that the system is located at a height of pc below the Galactic plane and has the size comparable to the scale height of the exponential cold gas disk (e.g., Velázquez and Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a). The interior of the nebula is also known to shine in soft thermal X-rays (Brinkmann et al., 1996, 2007; Safi-Harb and Ögelman, 1997; Chi et al., 2024), hard nonthermal X-rays (Seward et al., 1980; Watson et al., 1983; Yamauchi et al., 1994; Safi-Harb and Petre, 1999; Namiki et al., 2000; Safi-Harb et al., 2022; Kaaret et al., 2024), and filamentary H emission (Kirshner and Chevalier, 1980; Shuder et al., 1980; Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017; Rosado et al., 2021), which is indicative of the high-velocity shocks propagating through a relatively dense interstellar medium.
@@
The nature of the extended X-ray jets (EXJs) remains more puzzling, although their morphology, spectra, and polarization all point to the synchrotron origin of the emission (e.g., Safi-Harb et al., 2022; Kaaret et al., 2024). Most recently, very-high-energy (VHE) emission from these structures has been detected and mapped (e.g., Abeysekara et al., 2018; Alfaro et al., 2024; H. E. S. S. Collaboration et al., 2024; LHAASO Collaboration, 2025), demonstrating that ultra-relativistic electrons with energy above 100 teraelectronvolt (TeV) are accelerated in the system (e.g., Kimura et al., 2020), probably as a consequence of internal shocks and recollimation in the energetic trans-relativistic outflows (Bykov et al., 2025). The overall production of cosmic rays by such systems (super-Eddington accretors with powerful outflows) might be of importance for the total CR budget and local variations in its spectrum from place to place (e.g., Heinz and Sunyaev, 2002), particularly in the petaelectronvolt (PeV) regime (e.g., Peretti et al., 2025; Wang et al., 2025; Bykov et al., 2025).
@@
Although the paradigm of “a super-Eddington X-ray binary powering a nebula with energetic outflows” appears to be well established and consistent with the most salient observational properties of the SS 433/W50 complex, the details of almost all aspects of this interaction remain unclear. Indeed, all the phenomena directly associated with the central source can only be traced up to a distance of 0.1 pc from it, where the last signatures of the corkscrew pattern of the precessing jets are observed in radio and X-rays (Migliari et al., 2005; Miller-Jones et al., 2008; Khabibullin and Sazonov, 2017; Sakai et al., 2025). A model invoking steady deceleration and disruption of the jets can, in principle, account for such behavior (Panferov, 2014, 2017), but then the reappearance of the “jets” tens of parsec away from the central source with velocities similar to those at the injection might be problematic. Morphology of the soft X-ray emission, best mapped by the ROSAT mosaic observations (Brinkmann et al., 1996), is also quite puzzling, with large global brightness variations and thin filamentary structures, which are not commonly observed in similar situations of supernova remnants (SNRs, which typically feature shell-, plerion- or mixed types of morphology). Deep pointed observations with XMM-Newton, Chandra, Suzaku, and NuSTAR were capable of covering certain parts of the nebula (e.g., Brinkmann et al., 2005; Safi-Harb et al., 2022; Chi et al., 2024; Tsuji et al., 2025), but a uniform and sensitive coverage of the entire nebula with adequate spectroscopic capabilities is missing.
@@
Here, we present results of X-ray observations of the SS 433/W50 complex by the eROSITA telescope (Predehl et al., 2021) on board the SRG space observatory (Sunyaev et al., 2021), which were performed during its performance-verification (PV) phase. The obtained mosaic provides us with X-ray images of the entire extent of W50 with unprecedented dynamic range, bringing X-ray data on par with the radio and other multiwavelength data, and, in addition, delivers exquisite spectral information for the rich variety of diffuse structures. These data demonstrate a close connection between the thermal and nonthermal components of W50 on scales ranging from subparsec (thickness of narrow X-ray and radio bright filaments), to the entire pc extent of the nebula. These data also enable the characterization of the pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system (”extended X-ray jets”).
@@
The structure of the paper is as follows. We describe the observations and the obtained data in Sect. 2; results of the broad-band and spectrally resolved imaging are presented in Sect. 3 and Sect. 4, respectively; a spectroscopic analysis of the emission from several selected regions follows in Sect. 5; and the obtained global picture is discussed in Sect. 6 and summarized in Sect. 7.
@@ 2 Data
Since W50 spans approximately degrees on the sky, corresponding to a physical size of at a distance of 5 kpc111The most recent parallax measurement for SS 433 in Gaia DR3 corresponds to much larger distance of kpc (Gaia DR3 4293406612283985024, Gaia Collaboration et al., 2023); we advise caution if using this value given the large residuals of the astrometric solution. The previous parallax measurement from Gaia DR2 data gave a much smaller distance of kpc (Arnason et al., 2021)., the observing campaign consisted of nine individual telescope pointings (each covering a circular area of in diameter). These pointings were arranged in a mosaic to give a nearly uniform coverage of the entire nebula and, in addition, the highest sensitivity in the most interesting regions. The observing dates, from October 5 to October 8 and then from October 19 to October 20, 2019, were chosen to coincide with the two consecutive orbital eclipses of the central source SS 433 by its normal companion star (according to orbital ephemerids by Eikenberry et al., 2001). During eclipses, the apparent X-ray flux of SS 433 goes down by a factor of several (e.g., Kawai et al., 1989; Lopez et al., 2006; Medvedev et al., 2019), minimizing possible adverse effects of contamination and pile-up due to the presence of a very bright object in the field of view, even though the relative suppression of the soft X-ray lines was rather minor (Marshall et al., 2013). The typical effective (i.e., equivalent to the nominal eROSITA on-axis performance) exposure time across the map amounts to 30-50 ks, leading to images a factor of ten deeper compared to the previously available full-extent maps of W50 by the observatory in the 0.5-2 keV band (Brinkmann et al., 1996) and Einstein observatory in the 2-4 keV band (Watson et al., 1983). In addition to that, the eROSITA data have much better spatial ( on the telescope axis and averaged across the field of view) and spectral ( eV at 1 keV) resolution (Predehl et al., 2021). A comparable quality of the data is provided by multiple Chandra and XMM-Newton observations, but their coverage of the whole W50 nebula is nonuniform and still incomplete after more than 20 years of observations (Brinkmann et al., 2005; Safi-Harb et al., 2022; Chi et al., 2024). On the other hand, the SRG/eROSITA data are complementary to these observations, as well as to the imaging observations in harder X-rays with NuSTAR, which offer a unique view of nonthermally emitting populations. In this paper, we focus on the SRG/eROSITA X-ray data alone and leave combined analyses for future studies.
@@
We also supplement the PV data with the data obtained in the course of the SRG/eROSITA All-Sky Survey (Predehl et al., 2021; Sunyaev et al., 2021). Although the survey data are much shallower (the typical exposure time per point is s after combining the data of four scans), they provide uniform coverage of the entire area surrounding the nebula. For them, the standard data reduction, background estimation, and exposure and vignetting corrections were applied (similarly to mapping diffuse emission from nearby galaxy clusters (Churazov et al., 2021a, 2023), supernova remnants (Churazov et al., 2021b; Khabibullin et al., 2023, 2024), and the Galactic diffuse emission (Predehl et al., 2020; Khabibullin et al., 2022; Churazov et al., 2024) on even larger scales). Here, we also went one step further and combined the PV and all-sky data in a seamless fashion, similarly as it has been done for Chandra, XMM-Newton and SRG/eROSITA observations of the Perseus cluster (Churazov et al., 2025).
@@
A comparison of the degree fields in radio and X-ray bands is shown in Fig. 1. Although in the radio image the outer boundary of the W50 nebula appears to be very prominent, the X-ray image is instead dominated by the very bright central object (SS 433) and elongated structures known as extended X-ray jets (EXJs) located well inside the radio boundary of W50 (Seward et al., 1980; Watson et al., 1983; Brinkmann et al., 1996). At a distance of 5 kpc, these images span 350 pc and range from 17 pc to almost 370 pc in terms of distance from the Galactic plane. Given that the scale heights for the cold and warm components of the interstellar medium are tens and hundreds of pc, respectively, one could naturally expect signatures of the environment stratification (e.g., Goodall et al., 2011a). This is indeed revealed by the asymmetry of the W50 lobes, with the one directed toward the Galactic plane (the western lobe in equatorial coordinates) being more compact than the opposite one (the eastern lobe).
@@
On larger scales, clear variations of the background and foreground diffuse X-ray emission are seen. An inspection of 3D reddening maps (e.g., Green et al., 2019) showed a clear anticorrelation of the reddening and soft X-ray surface brightness (see Appendix C), suggesting that absorption of X-rays is the main reason for these variations. Morphologically similar structures appear in the 3D reddening maps at a distance of 2-3 kpc, i.e., in the foreground to W50/SS 433. Therefore, it is plausible that the majority of them are physically unrelated to W50.